OpenStudy (anonymous):

Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours. Use complete senten

OpenStudy (anonymous):

to compare, you must have the same units so TGV travels 89.44m.in a second = 89.44 * 3600 in an hour (one hour = 3600 seconds) = 321984 m in an hour = 32198/1000 = 321.98 km in an hour (one km = 1000 m) 321.98 > 320 so the TGV is faster than the Shinkansen Bullet Train 1 kilometer = 0.62137 miles so the TGV's speed is 321.98 * 0.62137 = 200.07 miles /hour and the distance traveled by the TGV in 2.5 hours is 200.07 * 2.5 = 500.17 miles. [;

OpenStudy (anonymous):

haha, thanks x

OpenStudy (anonymous):

Do i get a best response medal ? C: